OpenStudy (anonymous):

how do i simplify x^2-16/ x+4 ?

OpenStudy (mertsj):

\[\frac{x^2-16}{x+4}=\frac{(x+4)(x-4)}{x+4}=x-4\]

OpenStudy (anonymous):

how did you do that?

OpenStudy (mertsj):

What do you get when you factor x^2-16?

OpenStudy (anonymous):

you get 1,2,4,8

OpenStudy (mertsj):

x^2-16 is the difference of two squares and its factors are (x-4)(x+4)

OpenStudy (mertsj):

Then cancel the common factor x+4

OpenStudy (anonymous):

ok so then how would i solve \[y ^{2}+y-6/y ^{2}-4\] ? the factors of -6 are (x-2)(x+3)

OpenStudy (mertsj):

And the factors of \[y^2-4 are (y-2)(y+2)\]

OpenStudy (anonymous):

x+3/y+2?

OpenStudy (mertsj):

Cancel out the common factor y-2

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

ok so how would i find the restrictions?